### v1

@@ 3.1 Learning from Demonstrations
3.2 Deep Reinforcement Learning (DRL)
One area of robotics that has garnered interest [safe_learning] is Deep Reinforcement Learning (DRL), where Deep Neural Networks (DNN) predict the Q function of an MDP and the corresponding policy. It has been applied across tasks, e.g., grasping [QtOpt], locomotion [deeploco], and assembly [successfulAssembly].
Model-free, On-Policy DRL: Standard DRL gathers data from interaction with the environment, where Policy Gradient (PG) methods generate a batch of full trajectory data, compute the gradient of each state transition, and scale each gradient by the respective discounted return. PGs update the policy and repeat the process towards maximizing the expected cumulative reward. Proximal Policy Optimization (PPO) [PPO] is a PG method that leverages a value function, which is the average discounted return from a given state. These are on-policy methods, i.e., they only use data collected from the most recent policy. Exploration can be achieved through the periodic execution of random actions (epsilon greedy), injecting Gaussian noise to policy outputs [DDPG], or using a stochastic policy for entropy maximization [PPO].
DRL faces multiple challenges in robotics: (i) sample inefficiency: Policies need a lot of data to train and take significant training time. Real robot data, however, can be slow, expensive, and unsafe to collect [lessonslearnedtrainrobot]. (ii) instability: Policies can be inconsistent across training sessions due to poorly designed rewards, exploration strategies, learning rate parameters or learning error from DNNs. (iii) reward engineering: Simple rewards are sparse, i.e., they assign a 0 reward everywhere except at the goal. While desirable, sparse rewards often do not find a successful behavior in reasonable time [lessonslearnedtrainrobot]. This motivates dense rewards to guide exploration, which require manual engineering. While dense rewards can improve training time and stability, they can also lead to wrong behaviors, if they do not match the true task objective. (iv) long horizon tasks: The state space to be explored for such tasks is even larger making the propagation of rewards across subproblems difficult.
The above challenges have motivated multiple DRL variations in robotics:
@@ 3.2 Deep Reinforcement Learning (DRL)
Off-policy RL Sample inefficiency motivates data reuse. Off-policy RL stores the expected discounted future reward of any in-distribution state-action tuple on a Deep Q function, which is used to train the policy $\pi$. To update the parameters $\theta$ of the Deep Q function for a given state transition $(s,a,r,s^{\prime}$) (state, action, reward, next state), the immediate reward $r$ and future rewards $Q(s^{\prime},\pi_{\phi}(s^{\prime}))$ are used to define the loss function: $L(\theta)=E(Q_{\theta}(s,a)-r+\gamma\cdot Q_{\theta}(s^{\prime},\pi_{\phi}(s^{\prime})))^{2}$. This allows off-policy RL algorithms, such as TD3 [TD3] and SAC [SAC], to be more sample efficient than online RL. Off-policy RL, however, is complicated by the instability of DNN learning, especially because errors compound while learning Q functions [DQN]. A way to mitigate instability is to use target Q functions to query for future rewards [TD3][SAC][DQN]. The target Q functions average their weights slowly with the learned Q function’s weights. Given that the policy is defined as $\pi^{*}(s)=\arg\max_{a}Q(s,a)$, errors in Q function learning, which overvalue actions, are propagated rapidly. This is called an overestimation bias [TD3]. One way to mitigate overestimation bias to is use a clipped, double Q function, an ensemble of two Q networks coupled with target Q networks, that take the lower value of the two when queried for future rewards to underestimate Q values when uncertain [TD3][SAC]. Hindsight Experience Replay (HER) [HER] aims for sample-efficiency in multi-goal problems given only sparse rewards. It relabels the end or intermediate states of executed trajectories as the desired goal allowing to train over all available experiences.

### v2

@@ 3.1 Learning from Demonstrations
Figure 2: Robot motion generation methods that operate over an implicit model.
@@
3.2 Reinforcement Learning (RL)
One area of robotics that has garnered interest [safe_learning] is Reinforcement Learning (RL), where a machine learning model predicts the Q function of an MDP and the corresponding policy. It has been applied across tasks, e.g., grasping [QtOpt], locomotion [deeploco], and assembly [successfulAssembly].
Model-free, On-Policy RL: Standard RL gathers data from interaction with the environment, where Policy Gradient (PG) methods generate a batch of full trajectory data, compute the gradient of each state transition, and scale each gradient by the respective discounted return. PGs update the policy and repeat the process towards maximizing the expected cumulative reward. Proximal Policy Optimization (PPO) [PPO] is a PG method that leverages a value function, which is the average discounted return from a given state. These are on-policy methods, i.e., they only use data collected from the most recent policy. Exploration can be achieved through the periodic execution of random actions (epsilon greedy), injecting Gaussian noise to policy outputs [DDPG], or using a stochastic policy for entropy maximization [PPO].
RL faces multiple challenges in robotics: (i) sample inefficiency: Policies need a lot of data to train and take significant training time. Real robot data, however, can be slow, expensive, and unsafe to collect [lessonslearnedtrainrobot]. (ii) instability: Policies can be inconsistent across training sessions due to poorly designed rewards, exploration strategies, learning rate parameters or learning error from neural networks. (iii) reward engineering: Simple rewards are sparse, i.e., they assign a 0 reward everywhere except at the goal. While desirable, sparse rewards often do not find a successful behavior in reasonable time [lessonslearnedtrainrobot]. This motivates dense rewards to guide exploration, which require manual engineering. While dense rewards can improve training time and stability, they can also lead to wrong behaviors, if they do not match the true task objective. (iv) long horizon tasks: The state space to be explored for such tasks is even larger making the propagation of rewards across subproblems difficult.
The above challenges have motivated multiple RL variations in robotics:
@@ 3.2 Reinforcement Learning (RL)
Off-policy RL Sample inefficiency motivates data reuse. Off-policy RL stores the expected discounted future reward of any in-distribution state-action tuple on a Q function, which is used to train the policy $\pi$. To update the parameters $\theta$ of the Q function for a given state transition $(s,a,r,s^{\prime}$) (state, action, reward, next state), the immediate reward $r$ and future rewards $Q(s^{\prime},\pi_{\phi}(s^{\prime}))$ are used to define the loss function: $L(\theta)=E(Q_{\theta}(s,a)-r+\gamma\cdot Q_{\theta}(s^{\prime},\pi_{\phi}(s^{\prime})))^{2}$. This allows off-policy RL algorithms, such as TD3 [TD3] and SAC [SAC], to be more sample efficient than online RL. Off-policy RL, however, is complicated by the instability of neural network learning, especially because errors compound while learning Q functions [DQN]. A way to mitigate instability is to use target Q functions to query for future rewards [TD3][SAC][DQN]. The target Q functions average their weights slowly with the learned Q function’s weights. Given that the policy is defined as $\pi^{*}(s)=\arg\max_{a}Q(s,a)$, errors in Q function learning, which overvalue actions, are propagated rapidly. This is called an overestimation bias [TD3]. One way to mitigate overestimation bias to is use a clipped, double Q function, an ensemble of two Q networks coupled with target Q networks, that take the lower value of the two when queried for future rewards to underestimate Q values when uncertain [TD3][SAC]. Hindsight Experience Replay (HER) [HER] aims for sample-efficiency in multi-goal problems given only sparse rewards. It relabels the end or intermediate states of executed trajectories as the desired goal allowing to train over all available experiences.